# Diverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp-core enigma

### v1

@@ 2.2.2 CR modelling
Given that several of the CR physical processes are seemingly subdominant in the dwarf regime, Hopkins et al. (2023b) made a series of further simplifying assumptions that are applicable in a ‘streaming+diffusion’ limit. In particular, the reduced sub-grid model assumes that the CR energy equation is in steady state ($\partial_{t}e_{\mathrm{cr}}\to 0$) and that the magnetic fields below the resolution scale are isotropically ‘tangled’. The latter assumption allows the anisotropic diffusion tensor $\boldsymbol{\kappa}_{\mathrm{||}}$ to be replaced with an isotropic diffusion coefficient $\kappa$, averaged over the resolution scale (and spatio-temporally constant). The dominant loss term for CR energy is assumed to be a combination of hadronic/pionic, Coulomb, and ionization loses — therefore, neglecting losses from diffusive reacceleration (Hopkins et al., 2022b). Streaming contributions from advective and Alfvén velocities are also ignored, leading to a two parameter model depending on the isotropic diffusion rate $\kappa$ and a streaming velocity $v_{\mathrm{\kappa}}$. An approximate solution for the spatial distribution of CR energy density for such a model is presented in Hopkins et al. (2023b), and depends on an integral of an exponentially-declining function of the loss function.
@@ 3.1 Overview of the simulation suite
We begin our analysis of the simulation suite by inspecting visualisations of our simulated dwarfs’ stellar distributions. Figure 1 shows mock u/g/r Hubble images of the stars in each dwarf set-up, made by ray tracing through the gas: for each pixel, we sum up the luminosities of the star particles in each band along the line of sight. Dust associated with the gas elements attenuates the light, assuming a Milky Way-like attenuation curve in each band and a fixed dust-to-metals ratio. The side length of each projection is 10 kpc.
@@ 3.2 Stellar assembly
Figure 2: Integrated stellar properties of our simulations in comparison with observational constraints. The symbol style in this and later plots indicates the model variations. Left panel: Stellar mass – halo mass (SMHM) relation. SMHM relations based on empirical models as well as individual observed galaxies are plotted for comparison as indicated by the legend. Middle panel: (Intrinsic) simulated galaxy colours against stellar mass. The colours from the SDSS galaxies are indicated as grey contours and the integrated photometry of local dwarfs is plotted as grey squares (Mateo, 1998) and grey diamonds (Carlsten et al., 2022), for comparison. Right panel: Stellar half-mass radii against stellar mass with the data from local dwarfs (Carlsten et al., 2022) shown as grey diamonds for comparison. Overall, though the addition of AGN and/or CRs introduces significant scatter in the simulated distributions, our simulated stellar properties are in good agreement with the observations for all simulation set-ups explored.
Before assessing the impact of our baryonic modelling choices on dark matter distributions, it is vital to confirm that our feedback implementations produce realistic dwarf galaxy properties. To this end, we assess the stellar properties of our simulated dwarfs and compare these with observational constraints in Fig. 2, which shows the stellar mass – halo mass (SMHM) relation, galaxy colours and stellar half mass radii. Note that in this and later plots the symbol style indicates the model variations explored in our simulation suite, as indicated by the legend.
@@ 3.2.1 SMHM relation
For comparison, we plot several SMHM relations from the literature, including Behroozi et al. (2013, 2019) and Moster et al. (2018). The low-mass end of the SMHM relation remains largely unconstrained by observations, evident from the significant differences between these models. The Behroozi et al. (2013, 2019) models extend down to $\log(M_{\mathrm{halo}}/\mathrm{M_{\odot}})=10.0$, and the Moster et al. (2018) relation was shown to also be valid down to the dwarf regime, at least to $\log(M_{\mathrm{halo}}/\mathrm{M_{\odot}})=10.0$, by O’Leary et al. (2023). We plot the extrapolated relations by Behroozi et al. (2013, 2019) and Moster et al. (2018) as dashed and dotted lines, respectively. Extrapolating SMHM relations below the minimum mass considered has pitfalls, especially regarding physical processes like reionization suppression, which are effective at low masses but not included in most of the models. The scatter is especially difficult to model in this regime since more bursty star formation histories at lower stellar masses are expected to increase the scatter whilst high-redshift quenching due to reionization suppression may act to substantially reduce the scatter in the dwarf regime (O’Leary et al., 2023).
Our simulations with the m10q ICs, the least massive halo, are firmly in the extrapolated regime, whilst m10y and m10z are at the lower end of validity for the Behroozi et al. (2013, 2019) and Moster et al. (2018) relations. Hence for our simulations comparing to SMHM relations constructed for low-mass galaxy samples, rather than the fiducial SMHM relations for the general galaxy population, is more instructive. We include SMHM relations derived for dwarf galaxies by Read et al. (2017), Jethwa et al. (2018), and Nadler et al. (2020) as shaded regions. The differences in these relations arise due to different assumptions for surface-brightness incompleteness correction, see discussion in Behroozi et al. (2019). We also show various observed data points from the SPARC data base (Li et al., 2020) to indicate the scatter in the observations.
@@ 3.2.2 Galaxy colours
All simulated dwarfs are overall in good agreement with both the local dwarf galaxy constraints and the colour distribution of SDSS galaxies. Although we note that SDSS is very incomplete in the mass regime we consider here and the ELVES dwarfs were selected to be satellites while our dwarfs are field objects. The SDSS galaxies cluster around $g-r\sim 0.35$, whilst most of our dwarfs are offset towards somewhat redder colours, clustering around $g-r\sim 0.45$. Accounting for dust attenuation would likely make this offset more severe. However, this discrepancy is also inherently linked with completeness bounds in SDSS in this mass range, with blue galaxies likely overrepresented.
Interestingly, the m10qCR set-up has a similar stellar mass to its AGN counterpart m10qCR+AGN, but is shifted towards bluer colours due to a late-time burst in star formation which is also visible in the stellar projections in Figure 1. Conversely, the m10q set-up has a lower stellar mass and is redder than its AGN counterpart, indicating a complex interplay between CRs and AGN activity. For m10y, neither CRs nor the AGN have a strong impact on the integrated galaxy colours, though there are small, yet systematic, offsets with the CR runs being redder than the no CR runs, and the AGN runs being redder than their counterparts without AGN. Finally, for m10z, the set-up with neither CRs nor AGN yields the bluest galaxy colour, m10z, whilst the three other set-ups all have relatively similar colours, despite spanning an order of magnitude in stellar mass, indicating that all three of these formed the majority of their stars early in cosmic history.
@@ 3.2.3 Stellar sizes
We now turn towards examining the stellar sizes which provide crucial clues to both the star formation and dynamical histories of our dwarfs. We plot the stellar half mass radii of our simulated dwarfs at $z=0$ as colour-coded symbols. For comparison, we also plot the observed effective radii of local dwarf galaxies as grey diamonds based on the ELVES catalogue from Carlsten et al. (2022), which also includes the Milky Way dwarf measurements from McConnachie (2012) as a subsample. To make this comparison more consistent, we calculate the simulated stellar half mass radii based on the averages of 2D projections along random sightlines. Again there is appreciable scatter in the observations and our simulated dwarfs fall well within this scatter.
@@
Overall our stellar sizes are in good agreement with the observations, though the m10qCR+AGN simulation is slightly more compact than local dwarfs at similar stellar masses. However, we note that the observed radii are based on stellar light whilst we are calculating the simulated radii based on the stellar mass distribution. Observational effects can add significant scatter leading to both more compact half-light radii (e.g. Klein et al., 2024) or more extended observed distributions (e.g. Parsotan et al., 2021), with the stellar sizes also being significantly dependent on the wavelength (e.g. Cochrane et al., 2023).
@@ 3.3 BH assembly
We note that there is substantial uncertainty at the low-mass end of the BH – galaxy scaling relations due to the difficulty in reliably measuring the masses of intermediate-mass BHs (IMBHs) which have a much smaller gravitational sphere of influence and generally lower luminosities. Indeed there are no BH mass measurements for dwarf galaxies in the stellar mass range covered by our simulations, $10^{6}\ \mathrm{M_{\odot}}\lesssim M_{\mathrm{stellar}}\lesssim 10^{8}\ \mathrm{M_{\odot}}$. For more massive dwarfs there are only very limited measurements and we show these observed BH masses in massive dwarfs for reference, including virial BH mass estimates for active dwarf galaxies from Baldassare et al. (2020) and Wasleske & Baldassare (2024) as dark-grey squares and stars, respectively, and measurements from Greene et al. (2020) (based on dynamics, reverberation mapping or scaling from the stellar velocity dispersion in the specific case of TDE events that seem to have reliable determinations of the host galaxies) as dark-grey hexagons and triangles for mean values and upper limits, respectively. Error bars are omitted for clarity. We note that there are no dynamical nuclear IMBH mass measurements below $M_{\mathrm{stellar}}\sim 10^{9}\ \mathrm{M_{\odot}}$ and, excluding upper limits, the lowest BH mass estimate based on TDEs stems from a host galaxy of stellar mass $M_{\mathrm{stellar}}=2.5\times 10^{8}\ \mathrm{M_{\odot}}$. Below this mass, we are firmly in the extrapolated regime and indicate this by greying out the local BH – galaxy scaling relations below this mass.
@@ 3.4 Dark matter distributions
Fig. 4 shows the spherically averaged dark matter density profiles as a function of radius. The three panels display these profiles for our three sets of ICs, respectively. For these profiles, we centre the dark matter distributions on the minimum potential of the main halo [^3] Note that we also investigated alternative centering methods, including the shrinking sphere method from Power et al. (2003), however, we find that in practice the differences are minimal as the common halo centre definitions are in good agreement for our dwarfs.. In each case, we plot the mean $z=0$ profiles, averaged over $\Delta T=500$ Myr, for our four different galaxy formation model variations as colour-coded solid lines as indicated by the legend. We also performed additional dark-matter only (DMO) runs and we plot these as solid black lines. In each cases the temporal dispersion of these profiles is indicated by shaded regions which show the $1\sigma$ scatter in the densities over $\Delta T=500$ Myr. We note that for the majority of the simulation runs, the central density is not very sensitive to the temporal bin width as the evolution of the central density is relatively steady over the last Gyr. The m10zCR simulation, however, experiences significant fluctuations in the central density from 12similar-toabsent12$\sim$ 12∼ 1213131313 Gyr due to mergers with substructures, so for our $z=0$ profiles we choose the bin width to be small enough as to not be affected by these transient features. For all of the simulated profiles, we highlight the central dark matter density at $r=150$ pc, a characteristic radius for distinguishing cusps from cores in observations (Read et al., 2019), with colour-coded symbols.
Furthermore, we indicate the gravitational softening length at $r=62.5$ pc and the Power radius [^4] We note that the required particle number for resolving the central region depends on the radius and density contrast in question but generally ranges between $N=1000$–3000; here we follow Gutcke et al. (2021) and set $N=2000$. (enclosing the nearest 2000 neighbours, see Power et al., 2003) as grey and colour-coded dashed vertical lines, respectively. Within these radii, the dark matter profiles may be impacted by numerical heating effects. In addition, we show the expected NFW profile based on the virial mass as grey shaded regions indicating the $2\sigma$ scatter from the halo mass – concentration relation (Dutton & Macciò, 2014) to indicate the typical variations in profiles due to varying halo concentrations. The DMO runs are in good agreement with the NFW expectations, with m10y having slightly higher and m10z having slightly lower concentration than the mean relation.
@@
Figure 5: Central dark matter densities as a function of cosmic time. The dark matter densities are binned over $\Delta T=500$ Myr and the temporal $1\sigma$ scatter in each bin is shown by the shaded regions. The colour-coding indicates the respective simulation set-ups, as listed in the legend. For clarity, we plot the simulations without CRs as dotted lines. We measure the central densities at the inner radius $r=500$ pc to ensure that the central densities are well-resolved for all set-ups considered. For the m10q and m10y runs, the central densities are relatively steady at low redshifts, whilst the m10z runs still shows significant fluctuations due to a higher number of mergers and strong feedback events at late times.

### v2

@@ 2.2.2 CR modelling
Given that several of the CR physical processes are seemingly subdominant in the dwarf regime, Hopkins et al. (2023b) made a series of further simplifying assumptions that are applicable in a ‘streaming+diffusion’ limit. In particular, the reduced sub-grid model assumes that the CR energy equation is in steady state ($\partial_{t}e_{\mathrm{cr}}\to 0$) and that the magnetic fields below the resolution scale are isotropically ‘tangled’. The latter assumption allows the anisotropic diffusion tensor $\bm{\kappa}_{\mathrm{||}}$ to be replaced with an isotropic diffusion coefficient $\kappa$, averaged over the resolution scale (and spatio-temporally constant). The dominant loss term for CR energy is assumed to be a combination of hadronic/pionic, Coulomb, and ionization losses — therefore, neglecting losses from diffusive reacceleration (Hopkins et al., 2022b). Streaming contributions from advective and Alfvén velocities are also ignored, leading to a two parameter model depending on the isotropic diffusion rate $\kappa$ and a streaming velocity $v_{\mathrm{\kappa}}$. An approximate solution for the spatial distribution of CR energy density for such a model is presented in Hopkins et al. (2023b), and depends on an integral of an exponentially-declining function of the loss function.
@@ 3.1 Overview of the simulation suite
We begin our analysis of the simulation suite by inspecting visualisations of our simulated dwarfs’ stellar distributions. Figure 1 shows mock u/g/r Hubble images of the stars in each dwarf set-up, made by ray tracing through the gas: for each pixel, we sum up the luminosities of the star particles in each band along the line of sight. Dust associated with the gas elements attenuates the light, assuming a Milky Way-like attenuation curve in each band and a fixed dust-to-metals ratio. The side length of each projection is 10 kpc and the galaxies have been rotated to be viewed ‘face-on’ as defined by the angular momentum vector of the star particles.
@@
Figure 2: Star formation properties of our simulations. The symbol style in this and later plots indicates the model variations. Left panel: Median stellar age as a function of stellar mass with galaxy stellar ages from Local Group dwarfs shown as grey crosses for comparison (Weisz et al., 2014). Right panel: SFRs as a function of stellar mass. Observed SFRs for local star-forming dwarfs (Read & Erkal, 2019) and low-surface-brightness galaxies (McGaugh et al., 2017) are also shown for comparison as grey triangles. The inclusion of CRs leads to markedly older stellar populations and AGN-driven CR feedback results in quenched dwarfs at $z=0$ for all set-ups explored.
@@ 3.2 Stellar assembly
Before assessing the impact of our baryonic modelling choices on dark matter distributions, it is vital to confirm that our feedback implementations produce realistic dwarf galaxy properties. To this end, we assess the stellar properties of our simulated dwarfs in Figure 2. Note that in this and later plots the symbol style indicates the model variations explored in our simulation suite, as listed in the legend. We show the median stellar age as a function of stellar mass in the left panel. The galaxy stellar ages of Local Group dwarf galaxies from Weisz et al. (2014) as compiled by Mercado et al. (2021) are also plotted for comparison. In the right panel, we show the SFRs of our dwarfs at $z=0$ averaged over the past $500$ Myr as well as the observed SFRs of local dwarfs (Read & Erkal, 2019) and low-surface-brightness galaxies (McGaugh et al., 2017). The m10qCR and m10qCR+AGN simulations have markedly older stellar populations, corresponding to ancient ultra-faint dwarf galaxies (e.g. Brown et al., 2014), than their counterparts without CRs. The same trend holds for the other two dwarf haloes, albeit less pronounced. The SFRs of the star-forming simulated dwarf galaxies are broadly consistent with the star-forming dwarfs in the Local Group. Several of our dwarf simulations have been quenched for the past 500 Myr (indicated at the bottom of the right panel) consistent with the growing evidence for self-quenched dwarfs in low-density environments (Kaviraj et al., 2025). In particular, all simulations with AGN-driven CR feedback result in quenched dwarf galaxies at $z=0$. Finally, we note that the early formation of m10qCR and m10qCR+AGN also means that these set-ups are very sensitive to the assumptions on primordial chemical enrichment and, since we do not account for enrichment from Pop III stars, this leads to very low metallicities for these two simulated dwarfs ($\mathrm{[Fe/H]}\lesssim-3.0$, not shown), also see discussion in Wheeler et al. (2019).
Motivated, by this initial exploration of stellar assembly trends, we present further comparisons with observational constraints in Fig. 3, which shows the stellar mass – halo mass (SMHM) relation, galaxy colours and stellar half mass radii.
Figure 3: Integrated stellar properties of our simulations in comparison with observational constraints. Left panel: Stellar mass – halo mass (SMHM) relation. SMHM relations based on empirical models as well as individual observed galaxies are plotted for comparison as indicated by the legend. Middle panel: (Intrinsic) simulated galaxy colours against stellar mass. The colours from the SDSS galaxies are indicated as grey contours and the integrated photometry of local dwarfs is plotted as grey squares (Mateo, 1998) and grey diamonds (Carlsten et al., 2022), for comparison. Right panel: Stellar half-mass radii against stellar mass with the data from local dwarfs (Carlsten et al., 2022) shown as grey diamonds for comparison. Overall, though the addition of AGN and/or CRs introduces significant scatter in the simulated distributions, our simulated stellar properties are in good agreement with the observations for all simulation set-ups explored.
@@ 3.2.1 SMHM relation
For comparison, we plot several empirical SMHM relations from the literature, including Behroozi et al. (2013, 2019) and Moster et al. (2018). The low-mass end of the SMHM relation remains largely unconstrained by observations, evident from the significant differences between these models. The Behroozi et al. (2013, 2019) models extend down to $\log(M_{\mathrm{halo}}/\mathrm{M_{\odot}})=10.0$, and the Moster et al. (2018) relation was shown to also be valid down to the dwarf regime, at least to $\log(M_{\mathrm{halo}}/\mathrm{M_{\odot}})=10.0$, by O’Leary et al. (2023). Extrapolating SMHM relations below the minimum mass considered has pitfalls, especially regarding physical processes like reionization suppression, which are effective at low masses but not included in most of the models. The scatter is especially difficult to model in this regime since more bursty star formation histories at lower stellar masses are expected to increase the scatter whilst high-redshift quenching due to reionization suppression may act to substantially reduce the scatter in the dwarf regime (O’Leary et al., 2023). Indeed, there is also evidence from various studies that at very low halo masses ($\lesssim 10^{9}\ \mathrm{M_{\odot}}$) the SHMH relation may converge towards a constant stellar mass (e.g. Buck et al., 2019; Munshi et al., 2021; Arora et al., 2022; Monzon et al., 2024).
[…]
@@ 3.2.2 Galaxy colours
All simulated dwarfs are overall in good agreement with both the local dwarf galaxy constraints and the colour distribution of SDSS galaxies. Although we note that SDSS is very incomplete in the mass regime we consider here and the ELVES dwarfs were selected to be satellites while our dwarfs are field objects. The SDSS galaxies cluster around $g-r\sim 0.35$, whilst most of our dwarfs are offset towards somewhat redder colours, clustering around $g-r\sim 0.45$. Accounting for dust attenuation would likely make this offset more severe. However, this discrepancy is also inherently linked with completeness bounds in SDSS in this mass range, with blue galaxies likely overrepresented (Kaviraj et al., 2025).
Interestingly, the m10qCR set-up has a similar stellar mass to its AGN counterpart m10qCR+AGN, but is shifted towards bluer colours due to a late-time burst in star formation which is also visible in the stellar projections in Figure 1. Conversely, the m10q set-up has a lower stellar mass and is redder than its AGN counterpart, indicating a complex interplay between CRs and AGN activity. For m10y, neither CRs nor the AGN have a strong impact on the integrated galaxy colours, though there are small, yet systematic, offsets with the CR runs being redder than the no CR runs, and the AGN runs being redder than their counterparts without AGN. These trends can also be seen in the median stellar ages in Figure 2. Finally, for m10z, the set-up with neither CRs nor AGN yields the bluest galaxy colour, m10z, whilst the three other set-ups all have relatively similar colours, despite spanning an order of magnitude in stellar mass, indicating that all three of these formed the majority of their stars earlier in cosmic history as seen in the left-hand panel of Figure 2.
@@ 3.2.3 Stellar sizes
We now turn towards examining the stellar sizes which provide crucial clues to both the star formation and dynamical histories of our dwarfs. We plot the stellar half mass radii of our simulated dwarfs at $z=0$ as colour-coded symbols. For comparison, we also plot the observed effective radii of local dwarf galaxies as grey diamonds based on the ELVES catalogue from Carlsten et al. (2022), which also includes the Milky Way dwarf measurements from McConnachie (2012) as a subsample. To make this comparison more consistent, we calculate the simulated stellar half mass radii based on the averages of 2D projections along ten random sightlines with the components of the random direction vectors drawn from the normal distribution. Again there is appreciable scatter in the observations and our simulated dwarfs fall well within this scatter.
@@
Overall, our stellar sizes are in good agreement with the observations, though the m10qCR+AGNsimulation is slightly more compact than local dwarfs at similar stellar masses. However, we note that the observed radii are based on stellar light whilst we are calculating the simulated radii based on the stellar mass distribution. Observational effects can add significant scatter leading to both more compact half-light radii (e.g. Klein et al., 2024) or more extended observed distributions (e.g. Parsotan et al., 2021), with the stellar sizes also being significantly dependent on the wavelength (e.g. Cochrane et al., 2023).
@@ 3.3 BH assembly
We note that there is substantial uncertainty at the low-mass end of the BH – galaxy scaling relations due to the difficulty in reliably measuring the masses of intermediate-mass BHs (IMBHs) which have a much smaller gravitational sphere of influence and generally lower luminosities. Indeed there are no BH mass measurements for dwarf galaxies in the stellar mass range covered by our simulations, $10^{6}\ \mathrm{M_{\odot}}\lesssim M_{\mathrm{stellar}}\lesssim 10^{8}\ \mathrm{M_{\odot}}$. For more massive dwarfs there are only very limited measurements and we show these observed BH masses in massive dwarfs for reference, including virial BH mass estimates for active dwarf galaxies from Baldassare et al. (2020) and Wasleske & Baldassare (2024) as dark-grey squares and stars, respectively, and measurements from Greene et al. (2020) (based on dynamics, reverberation mapping or scaling from the stellar velocity dispersion in the specific case of tidal disruption events (TDEs) that seem to have reliable determinations of the host galaxies) as dark-grey hexagons and triangles for mean values and upper limits, respectively. Error bars are omitted for clarity. We note that there are no dynamical nuclear IMBH mass measurements below $M_{\mathrm{stellar}}\sim 10^{9}\ \mathrm{M_{\odot}}$ and, excluding upper limits, the lowest BH mass estimate based on TDEs stems from a host galaxy of stellar mass $M_{\mathrm{stellar}}=2.5\times 10^{8}\ \mathrm{M_{\odot}}$. Below this mass, we are firmly in the extrapolated regime and indicate this by greying out the local BH – galaxy scaling relations below this mass.
@@ 3.4 Dark matter distributions
Fig. 5 shows the spherically averaged dark matter density profiles as a function of radius. The three panels display these profiles for our three sets of ICs, respectively. For these profiles, we centre the dark matter distributions on the minimum potential of the main halo [^4] Note that we also investigated alternative centering methods, including the shrinking sphere method from Power et al. (2003), however, we find that in practice the differences are minimal as the common halo centre definitions are in good agreement for our dwarfs.. In each case, we plot the mean $z=0$ profiles, averaged over $\Delta T=500$ Myr, for our four different galaxy formation model variations as colour-coded solid lines as indicated by the legend. We also performed additional dark-matter only (DMO) runs and we plot these as solid black lines. In each cases the temporal dispersion of these profiles is indicated by shaded regions which show the $1\sigma$ scatter in the densities over $\Delta T=500$ Myr. We note that for the majority of the simulation runs, the central density is not very sensitive to the temporal bin width as the evolution of the central density is relatively steady over the last Gyr. The m10zCR simulation, however, experiences significant fluctuations in the central density from 12similar-toabsent12$\sim$ 12∼ 1213131313 Gyr due to mergers with substructures, so for our $z=0$ profiles we choose the bin width to be small enough as to not be affected by these transient features.
For all of the simulated profiles, we highlight the central dark matter density at $r=150$ pc, a characteristic radius for distinguishing cusps from cores in observations (Read et al., 2019), with colour-coded symbols. We caution when interpreting the densities at 150 pc that these are measured at a comparable scale to the gravitational softening of the dark matter and BH particles and therefore numerical effects may also impact the densities at these scales. For reference, we therefore indicate the gravitational softening length at $r\sim 89$ pc as a grey dashed line. To indicate the radii outside of which the impact of numerical heating should be minimised, we also show the Power radius [^5] We note that the required particle number for resolving the central region depends on the radius and density contrast in question but generally ranges between $N=1000$–3000; here we follow Gutcke et al. (2021) and set $N=2000$. (enclosing the nearest 2000 neighbours, see Power et al., 2003) for each simulation as colour-coded dashed vertical lines. We note, however, that strong temporal density fluctuations, indicative of numerical heating, are generally only seen at scales smaller than 150 pc so that we also include these scales in our analysis as a useful comparison point with observational work.
In addition, we show the expected NFW profile based on the virial mass as grey shaded regions indicating the $2\sigma$ scatter from the halo mass – concentration relation (Dutton & Macciò, 2014) to indicate the typical variations in profiles due to varying halo concentrations. The DMO runs are in good agreement with the NFW expectations, with m10y having slightly higher and m10z having slightly lower concentration than the mean relation.
@@
Figure 6: Central dark matter densities at $r=150$ pc (top panel) and $r=500$ pc (bottom panel) as a function of cosmic time. The dark matter densities are binned over $\Delta T=500$ Myr and the temporal $1\sigma$ scatter in each bin is shown by the shaded regions. The colour-coding indicates the respective simulation set-ups, as listed in the legend. For clarity, we plot the simulations without CRs as dotted lines. For the m10q and m10y runs, the central densities are relatively steady at low redshifts, whilst the m10z runs still shows significant fluctuations due to a higher number of mergers and strong feedback events at late times.